OpenStudy (anonymous):

A culture started with 6000 bacteria. After 5 hours, it grew to 7800 bacteria. Predict how many bacteria will be present after 13 hours. Round to the nearest whole number. (Exponential growth, P=Ae^kt)

OpenStudy (nurali):

7800=6000e^(5k) 1.3=e^(5k) ln (1.3)=5k k≈0.05 P=6000e^(0.05*13) P≈11493